# Topological phases, van Hove singularities, and spin texture in magic-angle twisted bilayer graphene in the presence of proximity-induced spin-orbit couplings

### v1

@@ I Introduction
Graphene materials, such as magic-angle twisted bilayer graphene (MATBG), bernal bilayer graphene, rhombohedral trilayer graphene, etc., possess strongly correlated and superconducting phases of matter [1, 2, 3, 4, 5, 6, 7, 8, 9, 3, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 15, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. Their properties can be tuned through accessible external parameters, such as gating, straining, and twist angles, providing valuable opportunities to study topology and correlated physics [33, 34, 35, 36]. In addition, experiments have utilized the proximity effect between graphene and a transition metal dichalcogenide (TMD) layer (such as WSe2), inducing proximity spin-orbit couplings (SOCs) in graphene and offering another method to control graphene-based materials [37, 38, 13, 26, 27, 28, 39, 40]. Interestingly, recent experiments have shown that observable superconductivity (SC) can be induced [13, 28, 41] or enhanced [26, 27] by the proximate $\text{WSe}_{2}$ layer. The interplay between SOC and SC in graphene is an active area of research [42, 43, 44, 45, 46, 47].
@@
Using the continuum Bistritzer-MacDonald (BM) model with Rashba and Ising SOCs [59], our findings show that the SOCs significantly reconstruct the band structure, splitting the two flat minibands (without SOC) into four spin-split bands. Each band has its own pair of van Hove singularities (VHSs), leading to a total of eight VHSs in DOS per valley, while only four VHSs per valley are present without SOC. Additionally, we discover a skyrmion-like spin texture in momentum space, which can evolve when crossing three distinct topological phases. Furthermore, we demonstrate that the spin texture in momentum space can be tuned in-plane, and the skyrmion-like feature can be further modified, by applying an out-of-plane electric field. Our results provide a systematic understanding of MATBG moiré bands in the presence of proximity-induced SOCs.
@@ II Model
We now address some subtleties concerning our model. First, the BM model is valid in the continuum limit and at low-energy and long wavelength. This is reasonable because, in this work, we only focus on the cases with isolated flat minibands with narrow bandwidth. Second, in our model, we use a simplified expression for the $\lambda_{R}$ term [64, 65, 68, 69]. The Rashba SOC will naturally be present when the mirror symmetry is broken, such that it can not only be induced by the proximate layer, but also an electric field perpendicular to the graphene sample. The proximity-induced SOC strengths depend on the relative angle between the SOC layer and the top graphene layer [70, 64, 71, 72, 68], while in the latter case, it depends on the strength of the electric field [57]. $\lambda_{I}$ and $\lambda_{R}$ can be as large as $\sim 20$ meV in DFT calculation[65], and $\sim 3$ meV from the tight-binding model [70], which is comparable to the bandwidth of MATBG. In general, it is well-known that an accurate knowledge of SOC from first principles calculations is a huge challenge, and it is more appropriate to obtain them by comparing with experiments. Here, we treat $\lambda_{I}$ and $\lambda_{R}$ as free parameters and study a range of the twist angles.
@@ III Single particle phase diagram, Berry phase and Chern number
Figure. 3 shows the contour plots of $\Delta_{23}(\textbf{k})$ and the Berry curvature of the second band $\Omega_{2}(\textbf{k})$ in the MBZ in k-space, near the A-B (panels a and b) and B-C (panels c and d) phase transitions, respectively. Around the A-B phase transition, there are three minima in the direct gap $\Delta_{23}$ located near $\Gamma_{M}$. When $E_{3}\left(\textbf{k}\right)-E_{2}\left(\textbf{k}\right)\rightarrow 0$ at a certain k point, Berry curvature tends to diverge at that point. Therefore $\Omega_{2}$ in the lower panel tends to diverge negatively in phase A, but positively in phase B, confirming the existence of a topological phase transition. On the other hand, near the B-C phase transition, there is only one minimum of $\Delta_{23}$ located directly at $\Gamma_{M}$, and $\Omega_{2}$ at $\Gamma_{M}$ tends to diverge positively in phase B and negatively in phase C. Generally speaking, the Berry curvature is extremely localized around the Dirac points near the phase boundaries.
@@ IV Density of States, Van Hove singularity
Now we discuss the impact of SOC to these VHSs. Without loss of generality, we take the case with $\lambda_{I}=\lambda_{R}=3$ meV and $\theta=1.05^{\circ}$ for an example. The situation in all three phases is similar. The proximity-induced SOC lifts the spin degeneracy, resulting in four spin-split bands: $E_{1},E_{2},E_{3},E_{4}$, as shown in Fig. 6a. The corresponding DOS are plotted in Fig. 6b and 6c. Interestingly, the VHS for each band splits into a pair of VHS, with energies very close to each other. These splittings result in a total of eight ordinary VHSs per valley, each exhibiting a logarithmically divergent DOS. The higher-order VHS observed in the absence of SOC disappears in the presence of SOCs. The precise VHS energies are mentioned in the caption of Fig. 6 and the corresponding energy contour plots are shown in Fig. 6d. Each VHS energy has three VHS points in k-space, where the Fermi pockets intersect at an angle. The VHS points are all close to the original Dirac points without SOC ($K_{M}$ and $K^{\prime}_{M}$). The splitting of the VHS is attributed to the breaking of the mirror symmetry around $k_{x}$ by SOC (but the C3 symmetry remains) Consequently, the energy $E_{k}$ is not the same at $K_{M}$ and $K^{\prime}_{M}$, causing the Fermi pockets to intersect at slightly different energies while remaining close to $K_{M}$ and $K^{\prime}_{M}$, unlike in Fig. 20. Therefore, for each pair of VHS, one is located near $K_{M}$ while the other is located near $K_{M}^{\prime}$.
@@ V.1 Emergent skyrmion-like spin texture in momentum space
We examine the spin texture of the four minibands for a range of twist angles, with $\lambda_{I}=\lambda_{R}=3$ meV. Figure. 8 presents the band structures for $1.03^{\circ}\leq\theta\leq 1.11^{\circ}$, where the color indicates the spin-$z$ expectation value: orange denotes spin-up, white denotes zero polarization, and purple denotes spin-down. For the angles presented, the system is in phase A, except for $1.08^{\circ}$, where the system is in phase C. Also, compared to $\theta=1.07^{\circ}$ and $1.09^{\circ}$, the cases with $\theta=1.03^{\circ}$ and $1.11^{\circ}$ are deeper in phase A because the energy difference between two middle bands at $\Gamma_{M}$, denoted $w_{0,\theta}$, is larger. Consequently, $\lambda_{I}/2w_{0,\theta}$ is smaller, indicating that these angles are further away from the A-B phase boundary.
@@
The spin texture profiles for $E_{1},E_{2},E_{3},E_{4}$, with $\theta=1.05$, are depicted in Figs. 9a-9d. The vector field represents the $S_{x}$ and $S_{y}$ components of the spin expectation value, and the spin texture adheres to $\mathcal{C}_{3}$ symmetry. For $E_{1}$, the spin vectors predominantly point downward (purple). At $\Gamma_{M}$, the spin vector points directly downward. Moving away from $\Gamma_{M}$, the spins tilt slightly away from the perpendicular direction, pointing away from $\Gamma_{M}$ in Zone 1 while remaining mostly downward in Zone 2. Zone 1 (pink) and Zone 2 (white) in the momentum space are defined on the left-hand side of Figs. 9a. Similarly, for $E_{4}$, the spins predominantly point upward (orange). At $\Gamma_{M}$, the spin vector points directly upward. In this case, the spins point toward $\Gamma_{M}$ in Zone 2 while remaining mostly upward in Zone 1. For $E_{2}$, the spin exhibit a skyrmion-like feature: the spin points purely upward at $\Gamma_{M}$ and downward at the corners of MBZ. Moving away from $\Gamma_{M}$, spins start to lie in-plane around the intersection of the orange and purple regions. Around $\Gamma_{M}$, the in-plane spin components point toward the center of the MBZ. On the other hand, for $E_{3}$, the trend is the opposite. At $\Gamma_{M}$, the spin points purely downward, while pointing upward at the corners of the MBZ, again indicative of a skyrmion-like feature. The spins also lie in-plane around the intersection of the orange and purple regions. However, in this scenario, the in-plane spin components around $\Gamma_{M}$ point away from the center of the MBZ. Moreover, on average, the in-plane spin components in Zone 1 are larger than that in Zone 2 for $E_{1},E_{3}$, and opposive for $E_{2},E_{4}$.
@@ V.2 Spin texture in presence of out-of-plane electric field
We now show how the spin texture evolves in the presence of an out-of-plane electric field. As mentioned before, the Rashba SOC is naturally present when the inversion symmetry is broken. Therefore, applying an out-of-plane electric field in MATBG also modifies the Rashba SOC. In this case, a dipolar coupling induces transitions between the $p_{z}$ and $s$ orbitals, flipping the spin [77, 78, 79, 57]. In MATBG, the out-of-plane electric field also gives rise to layer polarization, which can make the spin texture fully in-plane tunable [56]. Engineering SOC would influence the correlated phases [80, 81, 82, 83] and superconductivity [26, 84, 85, 42]. Specifically, engineering the Rashba SOC would control the polarization of spin accumulation and spin current [34, 56, 86]. In this section, we study how the in-plane spin components rotate around $\Gamma_{M}$ and $K_{M}$ points in the $+K$ valley. We find that around $K_{M}$, a radial Rashba spin texture can be readily achieved in MATBG, while around $\Gamma_{M}$, the spin texture can also be tuned to some extent in the presence of electric field.
@@
$h_{R}=\frac{\lambda_{R}}{2}e^{-i\psi s_{z}/2}\left(\tau\sigma_{x}s_{y}-\sigma_{y}s_{x}\right)e^{i\psi s_{z}/2}$
The Rashba angle actually depends on the twist angle between the upper layer of graphane and the SOC layer [65, 64, 50], which we ignored in the previous section. In this section, we show that the Rashba angle can be tuned from the radial ($\psi=90^{\circ}$) to the tangential ($\psi=0$) around the $K_{M}$ point by a displacement field.
@@
We present our results mainly at the twist angle $\theta=1.07^{\circ}$, because the effect of the out-of-plane electric field is more pronounced near the magic angle and right at magic angle, the skyrmion-like feature disappears (Fig. 8d). Later, we also comment on the results at other angles. As shown in Fig. 12, the bandwidth of the flat bands broadens as the displacement field increases. At large $u=80$ meV, the flat bands eventually merge into the higher energy bands, leading to the disappearance of the isolated flat bands (Fig. 12d). We limit our study to the vicinity of isolated flat bands, meaning $u<80$ meV. Overall, the top and bottom bands, $E_{1}$ and $E_{4}$, generally exhibit spin up/down characteristics even in the presence of an electric field. Interestingly, the spin-up/spin-down region for the $E_{2}$/$E_{3}$ in momentum space is significantly enhanced, as one increases the strength of the electric field: the position where $\langle S_{z}\rangle=0$ (white color) moves closer to $K_{M}$ instead of $\Gamma_{M}$ as $u$ increases.
@@ VI Discussion
We now mention several future directions. One can investigate possible impacts of the splitting of VHSs due to SOCs: Hall conductance, quantum oscillation in MATBG. In addition, one can also study the polarization of spin accumulation and spin current in the presence of SOC and out-of-plane electric field in MATBG, which should be able to serve as a platform with tailored electronic and spintronic properties. Inclusion of electron-electron interactions in the theory is also an important open and difficult challenge for the future.
@@ Appendix B Single-particle phase diagram with different interlayer hopping energy
In this Appendix, we use the interlayer hopping energy $w_{1}=110$ meV and $w_{0}/w_{1}=0.4$, to construct the single-particle phase diagram at twist angle $\theta=1.05^{\circ}$. As shown in Fig. 17, we still find three distinct topological phases A, B, and C. The A-B boundary is marked by purple, while the B-C boundary is marked by orange. The energy difference between the lower and upper bands at $\Gamma_{M}$ without SOCs, $w_{(0,\theta=1.05^{\circ})}=4.49$ meV here. We can see that with a smaller $w_{0}/w_{1}$ ratio (0.4), the width of phase B is smaller than that in the case with $w_{0}/w_{1}=0.8$ (Fig. 4e). In fact, in chiral limit, $w_{0}/w_{1}=0$, the phase B is extremely narrow. For example, when $\lambda_{R}=6$ meV, the width of phase B, $\lambda_{I}^{h}-\lambda_{I}^{l}\sim 1.403$ meV for $w_{0}/w_{1}=0.8$; $\sim 0.055$ meV for $w_{0}/w_{1}=0.4$; $\sim 0.006$ meV for $w_{0}/w_{1}=0$.

### v2

@@ I Introduction
Graphene materials, such as magic-angle twisted bilayer graphene (MATBG), bernal bilayer graphene, rhombohedral trilayer graphene, etc., possess strongly correlated and superconducting phases of matter [1, 2, 3, 4, 5, 6, 7, 8, 9, 3, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 15, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. Their properties can be tuned through accessible external parameters, such as gating, straining, and twist angles, providing valuable opportunities to study topology and correlated physics [34, 35, 36, 37, 38]. In addition, experiments have utilized the proximity effect between graphene and a transition metal dichalcogenide (TMD) layer (such as WSe2), inducing proximity spin-orbit couplings (SOCs) in graphene and offering another method to control graphene-based materials [39, 40, 13, 26, 27, 28, 41, 42, 43]. Interestingly, recent experiments have shown that observable superconductivity (SC) can be induced [13, 28, 44] or enhanced [26, 27] by the proximate $\text{WSe}_{2}$ layer. The interplay between SOC and SC in graphene is an active area of research [45, 46, 47, 48, 49, 50].
@@
Using the continuum Bistritzer-MacDonald (BM) model with Rashba and Ising SOCs [62, 63], our findings show that the SOCs significantly reconstruct the band structure, splitting the two flat minibands (without SOC) into four spin-split bands. Each band has its own pair of van Hove singularities (VHSs), leading to a total of eight VHSs in DOS per valley, while only four VHSs per valley are present without SOC. Additionally, we discover a skyrmion-like spin texture in momentum space, which can evolve when crossing three distinct topological phases. Furthermore, we demonstrate that the in-plane spin texture in momentum space can be tuned and, and the skyrmion-like feature can be further modified, by applying an out-of-plane electric field. Our results provide a systematic understanding of MATBG moiré bands in the presence of proximity-induced SOCs.
@@ II Model
We now address some subtleties concerning our model. First, the BM model is valid in the continuum limit and at low-energy and long wavelength. This is reasonable because, in this work, we only focus on the cases with isolated flat minibands with narrow bandwidth. Second, the large lattice constant mismatch between graphene (lattice constant 0.246 nm) and WSe2 (lattice constant 0.353 nm) result in a very weak moire effect (moiré lattice constant $\sim 1$ nm [13]), which we ignore. The proximity-induced SOC can have some spatial variations, but it averaging out quite rapidly within the TBG moire lattice constant $\sim 10$ nm. Third, in our model, we use a simplified expression for the $\lambda_{R}$ term [36, 68, 37, 59]. The Rashba SOC will naturally be present when the mirror symmetry is broken, such that it can not only be induced by the proximate layer, but also an electric field perpendicular to the graphene sample. The proximity-induced SOC strengths depend on the relative angle between the SOC layer and the top graphene layer [34, 36, 71, 72, 37], while in the latter case, it depends on the strength of the electric field [60]. $\lambda_{I}$ and $\lambda_{R}$ can be as large as $\sim 20$ meV in calculations and fitting to experiments [68, 34, 64, 73, 40, 74], which is comparable to the bandwidth of MATBG. In general, it is well-known that an accurate knowledge of SOC from first principles calculations is a huge challenge, and it is more appropriate to obtain them by comparing with experiments. Here, we treat $\lambda_{I}$ and $\lambda_{R}$ as free parameters and study a range of the twist angles.
@@ III Single particle phase diagram, Berry phase and Chern number
Figure. 3 shows the contour plots of $\Delta_{23}(\textbf{k})$ and the Berry curvature of the second band $\Omega_{2}(\textbf{k})$ in the MBZ, near the A-B (panels a and b) and B-C (panels c and d) phase transitions, respectively. Around the A-B phase transition, there are three minima in the direct gap $\Delta_{23}$ located near $\Gamma_{M}$. When $E_{3}\left(\textbf{k}\right)-E_{2}\left(\textbf{k}\right)\rightarrow 0$ at a certain k point, Berry curvature tends to diverge at that point. Therefore $\Omega_{2}$ in the lower panel tends to diverge negatively in phase A, but positively in phase B, confirming the existence of a topological phase transition. On the other hand, near the B-C phase transition, there is only one minimum of $\Delta_{23}$ located directly at $\Gamma_{M}$, and $\Omega_{2}$ at $\Gamma_{M}$ tends to diverge positively in phase B and negatively in phase C. Generally speaking, the Berry curvature, $\Omega_{2}$ and $\Omega_{3}$ are extremely localized around the Dirac points near the phase boundaries. On the other hand, the Berry curvature for the first and the fourth bands, $\Omega_{1}$ and $\Omega_{4}$, are much smaller in magnitude but still not flat enough to be considered similar to Landau levels, even though their variations are significantly less pronounced than in the second and third bands.
@@ IV Density of States, Van Hove singularity
Now we discuss the impact of SOC to these VHSs. Without loss of generality, we take the case with $\lambda_{I}=\lambda_{R}=3$ meV and $\theta=1.05^{\circ}$ for an example. The situation in all three phases is similar. The proximity-induced SOC lifts the spin degeneracy, resulting in four spin-split bands: $E_{1},E_{2},E_{3},E_{4}$, as shown in Fig. 6a. The corresponding DOS are plotted in Fig. 6b and 6c. Interestingly, the VHS for each band splits into a pair of VHS, with energies very close to each other. These splittings result in a total of eight ordinary VHSs per valley, each exhibiting a logarithmically divergent DOS. The higher-order VHS observed in the absence of SOC disappears in the presence of SOCs. The precise VHS energies are mentioned in the caption of Fig. 6 and the corresponding energy contour plots are shown in Fig. 6d. Each VHS energy has three VHS points in k-space, where the Fermi pockets intersect at an angle. The VHS points are all close to the original Dirac points without SOC ($K_{M}$ and $K^{\prime}_{M}$). The splitting of the VHS is attributed to the breaking of the mirror symmetry around $k_{x}$ by SOC (but the $\mathcal{C}_{3}$ symmetry remains). Consequently, the energy $E_{k}$ is not the same at $K_{M}$ and $K^{\prime}_{M}$, causing the Fermi pockets to intersect at slightly different energies while remaining close to $K_{M}$ and $K^{\prime}_{M}$, unlike in Fig. 20. Therefore, for each pair of VHS, one is located near $K_{M}$ while the other is located near $K_{M}^{\prime}$.
@@ V.1 Emergent skyrmion-like spin texture in momentum space
We examine the spin texture of the four minibands for a range of twist angles, with $\lambda_{I}=\lambda_{R}=3$ meV. Figure. 8 presents the band structures for $1.03^{\circ}\leq\theta\leq 1.11^{\circ}$, where the color indicates the spin-$z$ expectation value: orange denotes spin-up, white denotes zero polarization, and purple denotes spin-down. For the angles presented, the system is in phase A, except for $1.08^{\circ}$, where the system is in phase C. Also, compared to $\theta=1.07^{\circ}$ and $1.09^{\circ}$, the cases with $\theta=1.03^{\circ}$ and $1.11^{\circ}$ are deeper in phase A because the energy difference between two middle bands at $\Gamma_{M}$, denoted by $w_{0,\theta}$, is larger. Consequently, $\lambda_{I}/2w_{0,\theta}$ is smaller, indicating that these angles are further away from the A-B phase boundary.
@@
The spin texture profiles for $E_{1},E_{2},E_{3},E_{4}$, with $\theta=1.05^{\circ}$, are depicted in Figs. 9a-9d. The vector field represents the $S_{x}$ and $S_{y}$ components of the spin expectation value, and the spin texture adheres to $\mathcal{C}_{3}$ symmetry. For $E_{1}$, the spin vectors predominantly point downward (purple). At $\Gamma_{M}$, the spin vector points directly downward. Moving away from $\Gamma_{M}$, the spins tilt slightly away from the perpendicular direction, pointing away from $\Gamma_{M}$ in Zone 1 while remaining mostly downward in Zone 2. Zone 1 (pink) and Zone 2 (white) in the momentum space are defined on the left-hand side of Figs. 9a. Similarly, for $E_{4}$, the spins predominantly point upward (orange). At $\Gamma_{M}$, the spin vector points directly upward. In this case, the spins point toward $\Gamma_{M}$ in Zone 2 while remaining mostly upward in Zone 1. For $E_{2}$, the spins exhibit a skyrmion-like feature: the spin points purely upward at $\Gamma_{M}$ and downward at the corners of MBZ. Moving away from $\Gamma_{M}$, spins start to lie in-plane around the intersection of the orange and purple regions. Around $\Gamma_{M}$, the in-plane spin components point toward the center of the MBZ. On the other hand, for $E_{3}$, the trend is the opposite. At $\Gamma_{M}$, the spin points purely downward, while pointing upward at the corners of the MBZ, again indicative of a skyrmion-like feature. The spins also lie in-plane around the intersection of the orange and purple regions. However, in this scenario, the in-plane spin components around $\Gamma_{M}$ point away from the center of the MBZ. Moreover, on average, the in-plane spin components in Zone 1 are larger than that in Zone 2 for $E_{1},E_{3}$, and opposive for $E_{2},E_{4}$.
@@ V.2 Spin texture in presence of out-of-plane electric field
We now show how the spin texture evolves in the presence of an out-of-plane electric field. As mentioned before, the Rashba SOC is naturally present when the inversion symmetry is broken. Therefore, applying an out-of-plane electric field in MATBG also modifies the Rashba SOC. In this case, a dipolar coupling induces transitions between the $p_{z}$ and $s$ orbitals, flipping the spin [79, 80, 81, 60]. In MATBG, the out-of-plane electric field also gives rise to layer polarization, which can make the in-plane spin texture tunable [59]. Engineering SOC would influence the correlated phases [42, 82, 83, 84] and superconductivity [26, 27, 47, 45]. Specifically, engineering the Rashba SOC would control the polarization of spin accumulation and spin current [37, 59, 85]. In this section, we study how the in-plane spin components rotate around $\Gamma_{M}$ and $K_{M}$ points in the $+K$ valley. We find that, in the presence of electric field, a radial Rashba spin texture can be readily achieved around $K_{M}$, while around $\Gamma_{M}$, the spin texture can also be tuned to some extent in MATBG.
@@
$h_{R}=\frac{\lambda_{R}}{2}e^{-i\psi s_{z}/2}\left(\tau\sigma_{x}s_{y}-\sigma_{y}s_{x}\right)e^{i\psi s_{z}/2},$ (6)
which is originally derived for the induced Rashba SOC in monolayer graphene. In our study, we assume that the SOCs are induced by the proximity effect from the WSe2 layer, affecting only the top graphene layer. As a result, the Rashba SOC term follows the same form as in a monolayer due to the $\mathcal{C}_{3}$ symmetry. Although interlayer SOC terms involving layer pseudospin matrices could arise, they should be suppressed compared to the top-layer SOC terms in the proximity effect. This approach also aligns with experimental results, where SOC becomes significant only when carriers are polarized to the layer adjacent to the proximate TMD [26, 27, 44]. Futhermore, while the Rashba angle $\psi$ depends on the twist angle between the top graphene layer and the SOC layer [36, 53, 86, 68], we have neglected this effect in the previous section. Here, we set $\psi=0$ as before but show that the effective Rashba angle can be tuned from radial ($\psi=90^{\circ}$) to tangential ($\psi=0^{\circ}$) around the $K_{M}$ point by a displacement field.
@@
We present our results mainly at the twist angle $\theta=1.07^{\circ}$, because the effect of the out-of-plane electric field is more pronounced near the magic angle but right at magic angle, the skyrmion-like feature disappears (Fig. 8d). Later, we also comment on the results at other angles. As shown in Fig. 12, the bandwidth of the flat bands broadens as the displacement field increases. At large $u=80$ meV, the flat bands eventually merge into the higher energy bands, leading to the disappearance of the isolated flat bands (Fig. 12d). We limit our study to the vicinity of isolated flat bands, meaning $u<80$ meV. Overall, the top and bottom bands, $E_{1}$ and $E_{4}$, generally exhibit spin up/down characteristics even in the presence of an electric field. Interestingly, the spin-up/spin-down region for the $E_{2}$/$E_{3}$ in momentum space is significantly enhanced, as one increases the strength of the electric field: the position where $\langle S_{z}\rangle=0$ (white color) moves closer to $K_{M}$ instead of $\Gamma_{M}$ as $u$ increases.
@@ VI Discussion
We now mention several future directions. One is investigating the effects of VHS splitting due to SOC, such as its influence on Hall conductance and quantum oscillations in MATBG. Another potential direction is studying how a vertical electric field affects VHS splitting, as varying the electric field and SOC may lead to multiple transitions in the number of VHSs. In addition, one can also study the polarization of spin accumulation and spin current in the presence of SOC and out-of-plane electric field in MATBG, which should be able to serve as a platform with tailored electronic and spintronic properties. Inclusion of electron-electron interactions in the theory is also an important open and difficult challenge for the future.
@@ Appendix B Single-particle phase diagram with different interlayer hopping energy
In this Appendix, we use the interlayer hopping energy $w_{1}=110$ meV and $w_{0}/w_{1}=0.4$, to construct the single-particle phase diagram at twist angle $\theta=1.05^{\circ}$. As shown in Fig. 17, we still find three distinct topological phases A, B, and C. The A-B boundary is marked by purple, while the B-C boundary is marked by orange. The energy difference between the lower and upper bands at $\Gamma_{M}$ without SOCs, $w_{(0,\theta=1.05^{\circ})}=4.49$ meV here. We can see that with a smaller $w_{0}/w_{1}$ ratio (0.4), the width of phase B is smaller than that in the case with $w_{0}/w_{1}=0.8$ (Fig. 4e). In fact, in chiral limit, $w_{0}/w_{1}=0$, the phase B is extremely narrow. For example, when $\lambda_{R}=6$ meV, the width of phase B, $\lambda_{I}^{h}-\lambda_{I}^{l}\sim 1.403$ meV for $w_{0}/w_{1}=0.8$; $\sim 0.055$ meV for $w_{0}/w_{1}=0.4$; $\sim 0.005$ meV for $w_{0}/w_{1}=0$.